OpenStudy (anonymous):

Evaluate the following expression using the values given: Find x^3 - 2y^2 - 3x^3 + z^4 if x = 3, y = 5, and z = -3. ^=Exponent Please help soooo confused

OpenStudy (midhun.madhu1987):

Just plug in the value of x, y and z in the expression...

OpenStudy (anonymous):

3^3-2(5)^2-3(3)^3+(-3)^4= But how do i do all of this to get to the end result

OpenStudy (phi):

one term at a time (and be careful) what is 3^3 ? i.e. what is 3*3*3? write down the result then move on to the next term -2 * 5*5 = ? write it down can you do that ?

OpenStudy (anonymous):

3^3=9? -2 *5=10*5=50? Am i right or no?

OpenStudy (phi):

3^3 means 3*3*3 that is not 9

OpenStudy (midhun.madhu1987):

 similarly find each terms and find the value of the expression...